OpenStudy (anonymous):

Two functions, f(x) = 2x2 + 3 and g(x) = x + 1. If you are asked to find f[g(2)], which function would you solve first? Select one: a. f(x) b. g(x) c. Either f(x) or g(x) d. Not enough information given

OpenStudy (anonymous):

@sidsiddhartha @swift_13

OpenStudy (anonymous):

you solve g(x)

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

wait so the answer is B

OpenStudy (anonymous):

g(2)=3 and than plug 3 into f(x)

OpenStudy (anonymous):

is that the rightf(x) = 2x2 + 3 ?

OpenStudy (anonymous):

i think you tell me

OpenStudy (anonymous):

no i mean is that the right function you typed in? 2x2 (is that a x or multiplication sign?)

OpenStudy (anonymous):

its an x

OpenStudy (anonymous):

@sidsiddhartha

OpenStudy (anonymous):

well than it should be b. g(x)

OpenStudy (anonymous):

ok thank you soo much

OpenStudy (johnweldon1993):

So we're actually solving it now? well \[\large f(x) = 2x^2 + 3\] \[\large g(x) = x + 1\] f(g(2)) would mean to solve g(2) first...then lug the output in for 'x' in the 'f' function so \[\large g(2) = 2 + 1 = 3\] Now just plug in 3 for 'x' in the 'f' function \[\large f(g(2)) = 2(3)^2 + 3 = 2(9) + 3 = 18 + 3 = 21\]

OpenStudy (anonymous):

so is the correct answer is B right

OpenStudy (johnweldon1993):

..yes